Find the radius and centre of the circle:
(a) 3x^2 + 3y^2 + 12x + 18y - 43 = 0
1
Expert's answer
2021-01-26T04:22:06-0500

"3x^2 + 3y^2 + 12x + 18y - 43 = 0"

"3(x^2+4x+4)-12+3(y^2+6y+9)-27-43=0"

"3(x+2)^2+3(y+3)^2=82"

"(x+2)^2+(y+3)^2=\frac{82}{3}"

Answer: centre of circle is "(-2,-3)" and radius is "\sqrt{\frac{82}{3}}" .